OpenStudy (anonymous):

After solving the differential equation xy'=y-1, I end up with y-1 = x + C. How does this become y = Cx + 1?

OpenStudy (anonymous):

solving the DE gives ln(y-1)=ln(x)+k take k as lnC the u get y-1=Cx